# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### 18 Dec 2023 v2

@@ Appendix A The Influence of Noise on Task Entropy
The images $\boldsymbol{X}$ in the dataset are supposed to be independent of each other, as are the labels $\boldsymbol{Y}$. However, $\boldsymbol{X}$ and $\boldsymbol{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$, the conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{X}$ will depend on the joint distribution of $\boldsymbol{X}$ and $\boldsymbol{Y}$. Without knowing the joint distribution of $\boldsymbol{X}$ and $\boldsymbol{Y}$, we can not determine the conditional distribution of $\boldsymbol{Y}$ and $\boldsymbol{X}$. Here, we make some slacks for the distribution of $\boldsymbol{X}$ and $\boldsymbol{Y}$. We can transform the unknown distributions of $\boldsymbol{X}$ and $\boldsymbol{Y}$ to approximately conform to normality by utilizing some techniques, such as Box-Cox transformation, log transform, etc [5] [17]. After approximate transformation, the distribution of $\boldsymbol{X}$ and $\boldsymbol{Y}$ can be expressed as:
$\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu_{X}},\boldsymbol{\Sigma_{X}}),\boldsymbol{Y}\sim\mathcal{N}(\boldsymbol{\mu_{Y}},\boldsymbol{\Sigma_{Y}})$ (26)
@@
$\begin{split}&\boldsymbol{{\mu}_{X}}=\mathbb{E}[\boldsymbol{X}]=(\mathbb{E}[X_{1}],\mathbb{E}[X_{2}],...,\mathbb{E}[X_{k}]])^{T}\\ &\boldsymbol{{\mu}_{Y}}=\mathbb{E}[\boldsymbol{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\boldsymbol{\Sigma_{X}}=\mathbb{E}[(\boldsymbol{X}-\boldsymbol{{\mu}_{X}})(\boldsymbol{X}-\boldsymbol{{\mu}_{X}})^{T}]\\ &\boldsymbol{\Sigma_{Y}}=\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\end{split}$ (27)
@@
After transformation, the $\boldsymbol{X}$ and $\boldsymbol{Y}$ are subjected to multivairate normal distribution distribution. Then the conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{X}$ is also normally distributed [37] [23], which can be formulated as:
$\boldsymbol{Y}|\boldsymbol{X}\sim\mathcal{N}(\mathbb{E}({\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x}}),var(\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x}))$ (28)
where $\mathbb{E}({\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x}})$ is the mean of the label set $\boldsymbol{Y}$ given a sample $\boldsymbol{X}=\boldsymbol{x}$ from the dataset, and $var(\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x})$ is the variance of $\boldsymbol{Y}$ given a sample from the dataset. The conditional mean $\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x})]$ and conditional variance $var(\boldsymbol{Y}|\boldsymbol{X}=X)$ can be calculated as:
$\boldsymbol{\mu}_{\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x}}=\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x})]=\boldsymbol{{\mu}_{Y}}+\boldsymbol{\Sigma_{YX}}\boldsymbol{{\Sigma}^{-1}_{X}}(\boldsymbol{x}-\boldsymbol{\mu_{X}})$ (29)
$\boldsymbol{\Sigma}_{{\boldsymbol{Y|X}}=\boldsymbol{x}}=var(\boldsymbol{Y}|\boldsymbol{X}=\boldsymbol{x})=\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{X}\Sigma_{XY}}$ (30)
where $\boldsymbol{\Sigma_{YX}}$ and $\boldsymbol{\Sigma_{XY}}$ are the cross-covariance matrices between $\boldsymbol{Y}$ and $\boldsymbol{X}$, and between $\boldsymbol{X}$ and $\boldsymbol{Y}$, respectively, and $\boldsymbol{\Sigma^{-1}_{X}}$ denotes the inverse of the covariance matrix of $\boldsymbol{X}$.
Now, let $\boldsymbol{Z=Y|X}$, we shall obtain the task entropy:
$\begin{split}H(\mathcal{T};\boldsymbol{X})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X})\log{p(\boldsymbol{Y}|\boldsymbol{X})}\\ =&-\mathbb{E}[\log p(\boldsymbol{Y}|\boldsymbol{X})]\\ =&-\mathbb{E}[\log[(2\pi)^{-k/2}|\boldsymbol{\Sigma_{Z}}|^{-1/2}\exp(-\frac{1}{2}\boldsymbol{(Z-\mu_{Z})}^{T}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{Z}}(\boldsymbol{Z}-\boldsymbol{\mu_{Z}}))]]\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\boldsymbol{\Sigma_{Z}}|+\frac{1}{2}\mathbb{E}[{(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\boldsymbol{\Sigma_{Z}}|\end{split}$ (31)
@@
$\begin{split}\mathbb{E}[{(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})]=&\mathbb{E}[tr((\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}}))]\\ =&\mathbb{E}[tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}}){(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})}^{T})]\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}}){(\boldsymbol{Z}-\boldsymbol{\mu}_{\boldsymbol{Z}})}^{T})\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Z}}}\Sigma_{\boldsymbol{\boldsymbol{Z}}})\\ =&tr(\boldsymbol{I}_{k})\\ =&k\end{split}$ (32)
@@
Therefore, for a specific dataset, we can find that the task entropy is only related to the variance of the $\boldsymbol{Z}$.
@@
However, as we proactively inject additional information into the latent space, the task entropy changes and is defined as :
@@
$\left\{\begin{matrix}H(\mathcal{T};\boldsymbol{X}+\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})-H(\boldsymbol{X})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (35)
Eq. 35 differs from the conventional definition of conditional entropy, as our method injects the noise into the latent representations instead of the original images. If adding noise to the original images, then we have the classic definition:
$\left\{\begin{matrix}H(\mathcal{T};\boldsymbol{X}+\boldsymbol{\epsilon})=H(\boldsymbol{Y};\boldsymbol{X}+\boldsymbol{\epsilon})-H(\boldsymbol{X}+\boldsymbol{\epsilon})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\boldsymbol{X}\boldsymbol{\epsilon})=H(\boldsymbol{Y};\boldsymbol{X}\boldsymbol{\epsilon})-H(\boldsymbol{X}\boldsymbol{\epsilon})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (38)
Examples of the influence of various noises on the image level are provided in Fig. 3.
Figure 3: The influence of noise on the image. From left to right are the original image, the image with Gaussian noise, overlapping with its own linear transform, and with salt-and-pepper noise, separately.
A.1 Influence of Gaussian Noise on Task Entropy
Gaussian is one of the most common noises in image processing, and it is an additive noise. The Gaussian noise $\boldsymbol{\epsilon}$ is subjected to the normal distribution of $\boldsymbol{\epsilon}\sim\mathcal{N}(\mu_{\epsilon},\sigma_{\epsilon})$ and is independent of $\boldsymbol{X}$ and $\boldsymbol{Y}$. As we stated that the noise can be added to the original images or injected into the latent space, therefore, we discuss the conditions separately.
A.1.1 Add Gaussian Noise to Original Images
The task entropy with Gaussian noise is rewritten as:
$H(\mathcal{T};\boldsymbol{X+\epsilon})=-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X+\epsilon})\log{p(\boldsymbol{Y}|\boldsymbol{X+\epsilon})}$ (39)
Follow the derivations of the task entropy, we can calculate the task entropy with additive Gaussian noise as:
$\begin{split}H(\mathcal{T};\boldsymbol{X+\epsilon})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X+\epsilon})\log{p(\boldsymbol{Y}|\boldsymbol{X+\epsilon})}\\ =&-\mathbb{E}[\log p(\boldsymbol{Y}|\boldsymbol{X+\epsilon})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|X+\epsilon}}|\end{split}$ (40)
where $\boldsymbol{\Sigma_{Y|X+\epsilon}}=\boldsymbol{\Sigma_{Y}-\Sigma_{Y(X+\epsilon)}{\Sigma}^{-1}_{X+\epsilon}\Sigma_{(X+\epsilon)Y}}$. Since the Gaussian noise is independent of $\boldsymbol{X}$ and $\boldsymbol{Y}$, we have $\boldsymbol{\Sigma_{Y(X+\epsilon)}}=\boldsymbol{\Sigma_{(X+\epsilon)Y}}=\boldsymbol{\Sigma_{YX}}$. The corresponding proof is:
$\begin{split}\boldsymbol{\Sigma_{(X+\epsilon)Y}}=&\mathbb{E}\left[\boldsymbol{(X+\epsilon)}-\mu_{\boldsymbol{X+\epsilon}}\right]\mathbb{E}\left[\boldsymbol{Y}-\mu_{\boldsymbol{Y}}\right]\\ =&\mathbb{E}\left[\boldsymbol{(X+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(X+\epsilon)}\right]-\mu_{\boldsymbol{X+\epsilon}}\mathbb{E}\left[\boldsymbol{Y}\right]+\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{X+\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{(X+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(X+\epsilon)}\right]\\ =&\mathbb{E}\left[\boldsymbol{X}\boldsymbol{Y}\right]+\mathbb{E}\left[\boldsymbol{\epsilon}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{X}}-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{X}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{X}}\\ =&\boldsymbol{\Sigma_{XY}}\end{split}$ (41)
@@ A.1.1 Add Gaussian Noise to Original Images
Thus, the variation of task entropy adding Gaussian noise can be formulated as:
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Y|X}}|}{|\boldsymbol{\Sigma_{Y|X+\epsilon}}|}\leq 1\end{matrix}\right.$ (43)
To find the relationship between $|\boldsymbol{\Sigma_{Y|X+\epsilon}}|$ and $|\boldsymbol{\Sigma_{Y|X}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the data samples are independent of each other, and so are the labels.
@@
The same procedure can be applied to $\boldsymbol{\Sigma_{Y(X+\epsilon)}}$ and $\boldsymbol{{\Sigma}_{X+\epsilon}}$. Therefore, We can obtain that $\boldsymbol{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$, $\boldsymbol{\Sigma_{Y(X+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},X_{1}+\epsilon),...,\mathrm{cov}(Y_{k},X_{k}+\epsilon))$, and $\boldsymbol{{\Sigma}_{X+\epsilon}}$ is:
@@
$\begin{split}\boldsymbol{{\Sigma}_{X+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{X_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{X_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{X_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{X_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{X_{1}},...,\sigma^{2}_{X_{k}})\boldsymbol{I}_{k}+\sigma^{2}_{\epsilon}\boldsymbol{1}_{k}\end{split}$ (46)

### 02 Feb 2024 v3

@@ Appendix A Theoretical Foundations of Task Entropy
The image $\boldsymbol{X}$ in the dataset are supposed to be independent of each other, as are the labels $\boldsymbol{Y}$. However, $\boldsymbol{X}$ and $\boldsymbol{Y}$ are not independent because of the correlation between a data sample $X$ and its corresponding label $Y$. Essentially, the task entropy is the entropy of $p(Y|X)$. Following the principle of task entropy, compelling evidence suggests that diminishing task complexity via reducing information entropy can enhance overall model performance [30, 69].
Inspired by the concept of task entropy at the image level, we explore its extension to the latent space. The task entropy from the perspective of embeddings is defined as:
$H(\mathcal{T};\boldsymbol{Z})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})$ (32)
where $\boldsymbol{Z}$ are the embeddings of the images $\boldsymbol{X}$. Here, we assume that the embedding $\boldsymbol{Z}$ and the vectorized label $\boldsymbol{Y}$ follow a multivariate normal distribution. We can transform the unknown distributions of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ to approximately conform to normality by utilizing techniques such as Box-Cox transformation, log transform, etc., as described in [5] and [17] if they are not initially normal. After approximate transformation, the distribution of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ can be expressed as:
$\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\mu_{Z}},\boldsymbol{\Sigma_{Z}}),\boldsymbol{Y}\sim\mathcal{N}(\boldsymbol{\mu_{Y}},\boldsymbol{\Sigma_{Y}})$ (33)
@@
$\begin{split}&\boldsymbol{{\mu}_{Z}}=\mathbb{E}[\boldsymbol{Z}]=(\mathbb{E}[Z_{1}],\mathbb{E}[Z_{2}],...,\mathbb{E}[Z_{k}]])^{T}\\ &\boldsymbol{{\mu}_{Y}}=\mathbb{E}[\boldsymbol{Y}]=(\mathbb{E}[Y_{1}],\mathbb{E}[Y_{2}],...,\mathbb{E}[Y_{k}]])^{T}\\ &\boldsymbol{\Sigma_{Z}}=\mathbb{E}[(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})(\boldsymbol{Z}-\boldsymbol{{\mu}_{Z}})^{T}]\\ &\boldsymbol{\Sigma_{Y}}=\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\end{split}$ (34)
@@
Then the conditional distribution of $\boldsymbol{Y}$ given $\boldsymbol{Z}$ is also normally distributed [38] [24], which can be formulated as:
$\boldsymbol{Y}|\boldsymbol{Z}\sim\mathcal{N}(\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}={Z}}),var(\boldsymbol{Y}|\boldsymbol{Z}={Z}))$ (35)
where $\mathbb{E}({\boldsymbol{Y}|\boldsymbol{Z}=Z})$ is the mean of the label set $\boldsymbol{Y}$ given a sample $\boldsymbol{Z}=Z$ from the embeddings, and $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ is the variance of $\boldsymbol{Y}$ given a sample from the embeddings. The conditional mean $\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]$ and conditional variance $var(\boldsymbol{Y}|\boldsymbol{Z}=Z)$ can be calculated as:
$\boldsymbol{\mu}_{\boldsymbol{Y}|\boldsymbol{Z}=Z}=\mathbb{E}[(\boldsymbol{Y}|\boldsymbol{Z}=Z)]=\boldsymbol{{\mu}_{Y}}+\boldsymbol{\Sigma_{YZ}}\boldsymbol{{\Sigma}^{-1}_{Z}}(Z-\boldsymbol{\mu_{Z}})$ (36)
$\boldsymbol{\Sigma}_{{\boldsymbol{Y|Z}}=Z}=var(\boldsymbol{Y}|\boldsymbol{Z}=Z)=\boldsymbol{\Sigma_{Y}-\Sigma_{YX}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}$ (37)
where $\boldsymbol{\Sigma_{YZ}}$ and $\boldsymbol{\Sigma_{ZY}}$ are the cross-covariance matrices between $\boldsymbol{Y}$ and $\boldsymbol{Z}$, and between $\boldsymbol{Z}$ and $\boldsymbol{Y}$, respectively, and $\boldsymbol{\Sigma^{-1}_{Z}}$ denotes the inverse of the covariance matrix of $\boldsymbol{Z}$.
Now, we shall obtain the task entropy:
$\begin{split}H(\mathcal{T};\boldsymbol{Z})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{Z})\log{p(\boldsymbol{Y}|\boldsymbol{Z})}\\ =&-\mathbb{E}[\log p(\boldsymbol{Y}|\boldsymbol{Z})]\\ =&-\mathbb{E}[\log[(2\pi)^{-k/2}|\boldsymbol{\Sigma_{Z}}|^{-1/2}\exp(-\frac{1}{2}\boldsymbol{(Y|Z-\mu_{Y|Z})}^{T}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{Y|Z}}(\boldsymbol{Y|Z}-\boldsymbol{\mu_{Y|Z}}))]]\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|\end{split}$ (38)
@@
$\begin{split}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]=&\mathbb{E}[tr((\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}))]\\ =&\mathbb{E}[tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})]\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}\Sigma_{\boldsymbol{\boldsymbol{Y|Z}}})\\ =&tr(\boldsymbol{I}_{k})\\ =&k\end{split}$ (39)
@@
Therefore, for a specific set of embeddings, we can find that the task entropy is only related to the variance of the $\boldsymbol{Y|Z}$.
@@
As we proactively inject different kinds of noises into the latent space, the task entropy with noise injection is defined as :
@@
$\left\{\begin{matrix}H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\boldsymbol{Z}\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (42)
Equation 42 diverges from the conventional definition of conditional entropy as our method introduces noise into the latent representations instead of the original images. The noises examined in this study are classified into additive and multiplicative categories. In the subsequent sections, we analyze the changes in task entropy resulting from the injection of common noises into the embeddings.
Appendix B The Impact of Gaussian Noise on Task Entropy
We begin by examining the impact of Gaussian noise on task entropy from the perspective of latent space.
B.1 Inject Gaussian Noise in Latent Space
In this case, the task complexity is formulated as:
$\begin{matrix}H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})=H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})\end{matrix}.$ (43)
@@ B.1 Inject Gaussian Noise in Latent Space
Take advantage of the definition of task entropy, thus, the entropy change of injecting Gaussian noise in the latent space can be formulated as:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\mathcal{T};\boldsymbol{Z})-H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y}|\boldsymbol{Z})+H(\boldsymbol{Z})-(H(\boldsymbol{Y}|\boldsymbol{Z+\boldsymbol{\epsilon}})+H(\boldsymbol{Z+\boldsymbol{\epsilon}}))\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\log|\boldsymbol{\Sigma_{Z}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Z+\boldsymbol{\epsilon}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\end{split}$ (44)
where $\boldsymbol{\Sigma_{Y|Z+\epsilon}}=\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}$. Since the Gaussian noise is independent of $\boldsymbol{Z}$ and $\boldsymbol{Y}$, we have $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=\boldsymbol{\Sigma_{YZ}}$. The corresponding proof is:
$\begin{split}\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}-\mu_{\boldsymbol{Z+\epsilon}}\right]\mathbb{E}\left[\boldsymbol{Y}-\mu_{\boldsymbol{Y}}\right]\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]-\mu_{\boldsymbol{Z+\epsilon}}\mathbb{E}\left[\boldsymbol{Y}\right]+\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z+\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]+\mathbb{E}\left[\boldsymbol{\epsilon}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}\\ =&\boldsymbol{\Sigma_{ZY}}\end{split}$ (45)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}\leq 1\end{matrix}\right.$ (46)
To find the relationship between $|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|$ and $|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the embeddings of the images are independent of each other, and so are the labels.
@@
The same procedure can be applied to $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}$ and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$. Therefore, We can obtain that $\boldsymbol{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$,
$\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},Z_{1}+\epsilon),...,\mathrm{cov}(Y_{k},Z_{k}+\epsilon))$ (49)
and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ is:
@@
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{Z_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{Z_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}+\sigma^{2}_{\epsilon}\boldsymbol{1}_{k}\end{split}$ (50)